# Self-supervised prior learning improves structured illumination microscopy resolution

### v1

@@ Results
The ER forms a tubular network with 50–100 nm diameter tubules [Wang2022ERTubules]. Sparse-SIM provides sharper, more coherent results than the conventional SIM reconstruction in BioSR but fails to resolve tubular structures; Blind-SIM shows some microtubules but with many artifacts. SIMFormer clearly reveals tubules consistent with STORM results, and SIMFormer+ further improves robustness and suppresses low-light artifacts (Extended Data Fig. 4).
MTs are 25 nm-diameter cytoskeletal polymers essential for structural support and intracellular transport [Desai1997MTDynamics]. All methods resolve filamentous structures, but SIMFormer provides clearer results in overlapping regions, with SIMFormer+ further reducing artifacts. Extended Data Fig. 5 demonstrates that SIMFormer can resolve microtubules significantly narrower than the diffraction limit.
F-actin filaments (about 7 nm in diameter) form dense cytoskeletal networks [Cooper2022CellBook]. SIMFormer resolves these dense structures more clearly than other methods; SIMFormer+ further reduces artifacts in low-light regions (Supplementary Fig. 3). Resolution estimated by decorrelation analysis (Fig. 3e–h) shows that SIMFormer and SIMFormer+ reach average resolutions of 46.5 nm and 47.6 nm (4.9 and 4.8 times the diffraction limit), while conventional SIM reconstruction achieves around 1.7 times, Sparse-SIM about 2.3 times, and Blind-SIM about 3.1 times the diffraction limit but with significant artifacts [Descloux2019Decorrelation].
@@
The practical power of SIMFormer’s learned prior is especially evident in its ability to resolve $\sim 80$ nm ER tubules, a level of detail previously associated with STORM imaging [Wang2022ERTubules]. This is not merely image enhancement but reliable spectral extrapolation beyond the conventional $2\times$ resolution limit. The adaptive, data-driven prior overcomes the limitations of rigid handcrafted priors such as sparsity, which are ill-suited for complex, interconnected ER networks, and allows SIMFormer to infer high-frequency information inaccessible to traditional methods.
Light pattern prediction and noise robustness
@@
Trained by reconstructing raw SIM images, the accuracy of SIMFormer’s light pattern prediction is directly tied to the accuracy of its emitter predictions. We evaluate predicted light patterns on synthetic datasets and compare them to Blind-SIM. Due to learned priors and low-rank coding, SIMFormer’s predicted light patterns are significantly closer to the true ones, especially near emitters where light pattern accuracy matters most for reconstruction. On BioSR data, SIMFormer accurately predicts cosine-like light patterns across CCPs, ER, MTs, and F-actin, whereas Blind-SIM produces anomalous patterns (Fig. 4).
@@
Noise robustness is validated using both synthetic and BioSR datasets. With varying average photon numbers, SIMFormer and SIMFormer+ consistently outperform other methods in NRMSE and MS-SSIM. SIMFormer+ greatly improves SIMFormer’s performance under extremely low photon counts, reducing NRMSE and maintaining consistent resolution where SIMFormer alone suffers from artifacts that hinder decorrelation analysis (Fig. 5 and Supplementary Figs. 4–7).
On BioSR, side-by-side comparisons under low and high photon counts show that SIMFormer and SIMFormer+ produce fewer artifacts and greater robustness across tasks, with SIMFormer+ consistently improving artifact suppression and structure discernibility (Fig. 6 and Extended Data Figs. 6–9).
@@ Discussion
We have presented SIMFormer, a fully blind reconstruction framework that simultaneously estimates sample structure, PSF, light patterns, and background from raw SIM images. By combining a self-supervised learned prior with low-rank encoding of the light pattern, SIMFormer achieves a resolution approaching 45 nm—nearly five times the diffraction limit—while significantly improving light pattern estimation accuracy over previous blind methods [Mudry2012SpeckleSIM, Zhao2022SparseSIM]. With self-distillation, SIMFormer+ further enhances reconstruction accuracy and noise robustness without compromising resolution.
Conceptually, SIMFormer occupies an intermediate position between traditional MAP-based reconstruction and end-to-end supervised learning. Like MAP methods, it incorporates an explicit forward model of SIM imaging and uses this model to enforce consistency between predicted emitters, patterns, PSF, and raw data [Farsiu2004SR, Bouman1993GGMRF, Dey2006RLTV]. Unlike classical approaches, however, the regularizer is not a simple penalty such as sparsity or total variation; instead, it is an implicit prior encoded in the encoder–decoder network and learned directly from raw microscopy data via self-supervised training [He2022MAE, Feichtenhofer2022STMAE]. This design allows SIMFormer to adapt to complex biological structure statistics and to extrapolate beyond the nominal OTF cutoff, while maintaining robustness to experimental variability.
Compared with purely data-driven supervised approaches [Belthangady2019DLReview, Jin2020UNetSIM, Qiao2021DFCAN], SIMFormer avoids the need for large paired datasets of low- and high-resolution images, which are difficult or impossible to obtain for many biological preparations. Instead, it relies only on raw SIM stacks and a physically motivated forward model. This greatly simplifies data acquisition and makes the framework more amenable to deployment in diverse laboratories. Furthermore, because the forward model remains explicit, SIMFormer’s behavior is easier to interpret and diagnose than that of black-box image-to-image networks.
SIMFormer also demonstrates how learned priors and physical models can be tightly integrated. The architecture is deliberately structured so that each predicted component (emitters, patterns, background, PSF) has a clear physical interpretation, and each is regularized in a way that reflects prior knowledge: low-rank coding for patterns, smoothness and centering constraints for the PSF, and non-negativity for emitters and background. This modularity suggests clear paths for extension. For example, more elaborate PSF models could incorporate depth-dependent aberrations, and emitters could be modeled with additional constraints such as fluorophore blinking statistics or sparsity in specific transform domains.
The framework has several limitations. First, the current implementation is tailored to linear SIM and assumes a relatively simple cosine pattern model plus perturbations. Extending the method to non-linear SIM or more exotic illumination schemes will require careful redesign of the pattern branch. Second, while self-distillation significantly improves robustness at low photon counts, the training and fine-tuning procedures remain computationally demanding, relying on modern GPUs or TPUs [Bradbury2018JAX]. Third, although we observe good generalization across structures and imaging systems, a systematic study of failure modes—particularly in samples that strongly violate the statistics of the training data—will be important for future work.
Beyond SIM, the ideas underlying SIMFormer are applicable to a broad class of inverse problems in optical imaging, where forward models are well understood but reconstruction is ill-posed. By combining flexible learned priors with explicit physical models and self-supervision, similar architectures could be developed for applications such as light-sheet microscopy, adaptive optics, or computational phase imaging [Tian2025CompSuperRes, Wang2024ImagingMechanism]. Finally, the relationship between SIMFormer’s reconstruction error and information-theoretic bounds, such as the Cramér–Rao lower bound, remains an open question. Analyzing this relationship would clarify how close data-driven priors can bring practical reconstruction performance to fundamental physical limits.
@@
As schematically illustrated in Fig. 1 and Extended Data Fig. 1, the encoder is designed to process raw SIM images and is inspired by spatiotemporal masked autoencoders [Feichtenhofer2022STMAE]. A ViT-based encoder with pre-trained CLIP weights [Dosovitskiy2021ViT, Radford2021CLIP] processes multi-frame SIM images. To speed up training and improve extraction of fine-grained features, we incorporate a ViT-adapter [Chen2023ViTAdapter].
Multi-frame SIM images are converted into multiple 768-dimensional vectors via 3D patch embedding. In parallel, the original images are processed through a CNN-based stem, yielding an equal number of 768-dimensional vectors per patch. Both sets are subjected to the same random masking. Unmasked patches are combined with learnable position and frame embeddings, shuffled, and fed into encoder and adapter layers, enabling interaction and information exchange. The encoder output is projected to lower-dimensional vectors per patch; a learnable token replaces masked patches. These features are reordered, combined with positional embeddings, and fed into a decoder consisting of multiple Transformer layers, whose outputs feed three prediction heads.
@@
Emitters and background are predicted by a CNN-based module with four blocks, each performing two-fold upsampling. Within each block, upsampling is followed by three $3\times 3$ convolutions with Gaussian error linear unit (GELU) activations. The module starts with 512 channels and halves the channel count at each block; a final $1\times 1$ convolution produces a single-channel output.
The emitter prediction branch consists of this module followed by a Softplus activation to enforce non-negative outputs. The background branch uses the same module followed by an average pooling layer with kernel size 64, reflecting the assumption that the background varies slowly in space.
@@
In SIM, the light pattern consists of only a few modes. A typical SIM experiment uses cosine patterns in three directions with three phases per direction, corresponding to nine distinct modes [Gustafsson2000SIM, Mudry2012SpeckleSIM]. To encode this prior knowledge, we use low-rank encoding via average pooling. Decoder features are first passed through a Transformer layer, then through one-dimensional average pooling applied to each feature vector, and finally through another Transformer layer to produce the light pattern output.
A common assumption is that the sum of all light patterns in a SIM experiment yields uniform illumination [Mudry2012SpeckleSIM]. To ensure that the sum equals one and intensities remain non-negative, we apply a softmax operation along the channel dimension. During multi-stage training, the effective rank of the pattern representation is gradually reduced, biasing the solution towards simpler, more interpretable patterns while preserving the ability to capture experimental deviations from ideal cosines.
@@
The PSF prediction module is a CNN that maps a fixed random tensor to the PSF. Following the idea of Deep Image Prior [Ulyanov2018DeepImagePrior], the network recasts PSF estimation as an optimization problem within a neural network, and has been shown to learn PSFs directly from raw microscopy images [Wang2024ImagingMechanism]. We incorporate total-variation regularization on the PSF to encourage smoothness and penalize unphysical oscillations. A centering loss aligns the intensity-weighted PSF centroid with the geometric center of the tensor, preventing drift during training.
@@
The loss function combines several terms to enable self-supervised learning from raw SIM images. The first category measures differences between reconstructed and raw SIM images, using an $L_{1}$ distance and MS-SSIM; their weighted average is known to be effective for image restoration [Zhao2016LossFunctions]. The second category includes: (1) a center-distance term enforcing alignment between the intensity-weighted PSF center and tensor center; and (2) total-variation losses on PSF and light patterns to encourage spatial smoothness. The overall loss is a weighted sum of these terms, with hyperparameters chosen to balance data fidelity and regularization across synthetic and real datasets.
Training procedure
Training begins from a pre-trained CLIP-based initialization [Radford2021CLIP], since training ViTs from scratch on small datasets is suboptimal [He2022MAE]. Initial training uses batch size 96, 70 epochs, patch size $3\times 16\times 16$, and mask ratio 0.75. A multi-stage schedule then gradually compresses the light pattern rank from 512 to 32 (256, 128, 64, 32), with each stage using mask ratio 0.25, batch size 64, and 30 epochs. After reaching rank 32, we train for 2000 epochs with batch size 96 and mask ratio 0.25 for synthetic data / 0.75 for BioSR.
Throughout, patch size $3\times 16\times 16$, crop size $80\times 80$, PSF size $49\times 49$, and learning rate $10^{-4}$ are maintained. Raw images are linearly scaled by a factor of three, yielding threefold-enlarged SIM reconstructions; unscaled raw images are used in the loss to maintain physical consistency with the forward model.
@@
To obtain SIMFormer+, we perform self-distillation. SIMFormer is first trained on high-SNR SIM images, producing high-resolution predictions. These are then used as soft labels to train SIMFormer+ under lower photon counts and higher noise. High-SNR versions of test images are excluded from both training sets to avoid information leakage. SIMFormer+ thus inherits high-resolution capabilities while greatly improving noise robustness [Zhang2019SelfDistill, Zhang2022SelfDistillationSurvey, Pham2022RevisitingSelfDistill]. In practice, we observe that SIMFormer+ suppresses spurious high-frequency artifacts that sometimes appear in the teacher’s predictions when photon counts are extremely low.
@@
We apply percentile-based normalization to both ground truth and reconstructed images:
$I_{\mathrm{norm}}=\frac{I-\mathrm{Percentile}(I,0.1)}{\mathrm{Percentile}(I,99.9)-\mathrm{Percentile}(I,0.1)}.$
This ensures consistent dynamic ranges across images and reduces sensitivity to outliers and hot pixels. All metrics (NRMSE, MS-SSIM) are computed on normalized images.
@@
Resolution is assessed via decorrelation analysis [Descloux2019Decorrelation], which identifies the highest spatial frequency by examining local peaks in decorrelation functions, rather than relying on Abbe’s criterion. Briefly, the decorrelation function is computed in the Fourier domain and radial profiles are extracted. Local maxima in these profiles correspond to effective resolution limits. We follow the procedure described in Ref. [Descloux2019Decorrelation] and apply it uniformly to all reconstruction methods for fair comparison.
Computational resources and speed
We implement SIMFormer in JAX [Bradbury2018JAX], training on eight NVIDIA Tesla A100 80 GB GPUs for seven days (24 hours/day). Inference for a single SIM stack of shape $(9,512,512)$ takes approximately one second. All experiments are performed using mixed-precision training to reduce memory consumption and accelerate convergence.

### v2

@@ Results
The ER is a key organelle involved in protein synthesis, lipid metabolism, and calcium storage, forming a tubular network with nanoscale tubules around 50–100 nm in diameter, as confirmed by STORM microscopy [Wang2022ERTubules]. These tubules play a critical role in material distribution and maintaining cellular homeostasis. Fig. 3b presents a side-by-side comparison of ER images, showing that while Sparse-SIM provides sharper and more coherent results than BioSR SIM, it fails to resolve the tubular structures. Blind-SIM shows some microtubules but introduces many artifacts. In contrast, SIMFormer clearly reveals the 80 nm width microtubules closely matching STORM results. SIMFormer+ achieves similar results while improving robustness and suppressing artifacts in low-light regions. More results are shown in Extended Data Fig. 4.
MTs are essential components of the cytoskeleton, providing structural support, facilitating intracellular transport, and playing key roles in cell division and signaling. These dynamic, tube-like polymers, typically 25 nm in diameter, are composed of tubulin proteins and are critical for maintaining cell shape and organizing organelles [Desai1997MTDynamics]. Fig. 3c shows the comparison results for MTs, with all methods resolving the filamentous structures. However, SIMFormer offers clearer results in regions where the microtubules overlap, particularly in the zoomed-in sections highlighting intersecting structures. Additionally, SIMFormer+ further reduces artifacts. Extended Data Fig. 5 demonstrates that SIMFormer can resolve microtubule structures as narrow as 34 nm in width.
F-actin (filamentous actin) is a critical component of the cytoskeleton, playing a vital role in cellular processes such as movement, shape maintenance, and division. These thin, flexible filaments, about 7 nm in diameter, are made up of actin subunits and form dense networks that provide structural support and facilitate cellular motility [Cooper2022CellBook]. Fig. 3d presents the comparison results for F-actin filaments. SIMFormer resolves these dense filamentous structures more clearly than other methods, while SIMFormer+ further reduces artifacts in low-light regions. Additional results are shown in Supplementary Fig. 3.
Fig. 3e–h shows the resolution calculated using decorrelation analysis for BioSR SIM, Sparse-SIM, Blind-SIM, SIMFormer, and SIMFormer+ across the ER, CCPs, MTs, and F-actin filaments datasets. SIMFormer and SIMFormer+ demonstrate superior resolution compared to the other methods, reaching on average 46.5 nm and 47.6 nm (4.9 and 4.8 times the diffraction limit), consistent with their previously observed resolving capabilities in the synthetic data experiment. SIMFormer+ shows minimal loss in resolution while significantly improving robustness and suppressing artifacts. In contrast, conventional SIM reconstruction (BioSR SIM) achieves 1.7 times the diffraction limit, Sparse-SIM reaches 2.3 times, and Blind-SIM, while achieving 3.1 times the diffraction limit, suffers from significant artifacts.
@@
The practical power of SIMFormer’s learned prior is clearly demonstrated by its ability to resolve 80 nm ER tubules, a level of detail previously associated with STORM imaging [Wang2022ERTubules]. This achievement is not merely image enhancement; it represents a form of reliable spectral extrapolation beyond the conventional 2x resolution limit. Critically, the learned prior overcomes the limitations of hand-crafted rules, such as sparsity, which are ill-suited for complex biological structures like the interconnected ER network. Reliance on such rigid priors can cause algorithms like Sparse-SIM to misinterpret these dense tubular structures as simple fine lines. In contrast, SIMFormer’s adaptive, data-driven prior allows it to plausibly infer high-frequency information that is inaccessible to traditional methods, thereby breaking the conventional resolution barrier to reveal these ultrastructures with higher fidelity.
SIMFormer predicts accurate light patterns with self-supervised learned priors
@@
Trained by reconstructing raw SIM images, the accuracy of SIMFormer’s light pattern prediction is directly tied to the accuracy of its emitter predictions. We evaluate SIMFormer’s light pattern prediction on synthetic datasets and compare it to Blind-SIM. As shown in Fig. 4d, SIMFormer achieves an average NRMSE of 0.39, while Blind-SIM records 0.51. Notably, the light pattern evaluation is performed only in regions near the emitters (Supplementary Note 2), as the reconstruction of the light pattern in areas far from the emitters has no impact on the emitter reconstruction. As shown in Fig. 4a–c, due to the learned priors and low-rank coding, SIMFormer’s predicted light patterns are significantly closer to the true ones, supporting the quantitative comparison. Fig. 4e–h shows light pattern predictions for CCPs, ERs, MTs, and F-actin from the BioSR dataset. SIMFormer accurately predicts the cosine light pattern across all structures, whereas Blind-SIM produces anomalous light patterns.
SIMFormer achieves both high resolution and strong noise robustness through self-distillation
@@
We validate the noise robustness of SIMFormer and SIMFormer+ using both synthetic and BioSR datasets. In the synthetic datasets, we test the three previously mentioned structures, with a light pattern spatial period of 313 nm and average photon numbers set to 0.5, 1, 10, 100, 500, and 1000, respectively. The quantitative results are shown in Fig. 5b. NRMSE and MS-SSIM measurements across varying photon numbers indicate that SIMFormer and SIMFormer+ comprehensively outperform all other methods. A detailed analysis of the NRMSE values reveals that SIMFormer+ significantly enhances SIMFormer’s performance under extremely low photon counts (0.5), with a 30.1% reduction in NRMSE for synthetic filament structures, 4.8% for synthetic microtubule structures, and 17.5% for synthetic ring structures. As shown in Supplementary Fig. 4, SIMFormer+ maintains consistent resolution at low photon counts, while SIMFormer struggles due to artifacts that prevent decorrelation analysis. The observations in Fig. 5b further support these findings, demonstrating that SIMFormer+ greatly improves SIMFormer’s robustness under low photon counts. Additional results are provided in Supplementary Figs. 5–7.
In Fig. 6, we present a side-by-side comparison of conventional SIM reconstruction, SIMFormer, and SIMFormer+ on the BioSR dataset under both low and high photon counts. To ensure that SIMFormer does not memorize low-noise data during the noise robustness evaluation, we use the first 35 cells as the training set and the remaining cells as the test set. Key conclusions are as follows: (1) SIMFormer and SIMFormer+ consistently produce fewer artifacts and greater robustness under low photon counts across all tasks, (2) SIMFormer+ generates fewer artifacts than SIMFormer under low photon counts, and (3) SIMFormer+ enhances the discernibility of biological structures, such as the ring-shaped structures in CCPs and the tubular structures in ER, compared to SIMFormer under low photon counts. Further results can be found in Extended Data Figs. 6–9. These findings suggest that the learnable priors in SIMFormer, combined with the self-distillation in SIMFormer+, deliver a significant improvement in both resolution and noise robustness.
@@ Discussion
In this work, we present SIMFormer, a fully blind reconstruction framework that simultaneously estimates the sample structure, PSF, light patterns, and background from raw SIM images. By integrating a learnable prior acquired through self-supervised learning with a low-rank encoding of the light pattern, SIMFormer achieves a resolution approaching 45 nm—nearly five times the diffraction limit—while significantly improving the accuracy of light pattern estimation over previous blind methods. Furthermore, with the introduction of a self-distillation step, the resulting SIMFormer+ model enhances both reconstruction accuracy and noise robustness without compromising this high resolution.
The remarkable performance of SIMFormer stems from its novel approach to learning physics-informed priors. Instead of relying on handcrafted rules, the model’s architecture is designed to follow SIM imaging principles, while its self-supervised training on raw data compels it to capture the intrinsic statistical distribution of complex biological structures. This synergy creates a powerful, data-driven regularizer that effectively constrains the solution space for the ill-posed inverse problem, thereby reducing reconstruction ambiguity. This learned prior is also modular and could be combined with traditional handcrafted priors, such as sparsity or continuity, to potentially achieve further performance gains.
SIMFormer’s design principles also naturally suggest routes for extending data-driven priors to more complex imaging scenarios, without relying on additional calibration data. By tightly coupling physical image formation models with flexible learned priors, the framework provides a template for future algorithms that must balance robustness, fidelity, and practicality in demanding experimental settings.
Furthermore, the framework can be naturally adapted to process volumetric data by using 3D patches and adding axial positional encodings. Such an extension holds the promise of achieving isotropic 3D super-resolution, which, combined with the method’s inherent noise robustness, would be particularly powerful for gentle, long-term 4D live-cell imaging of dynamic processes without causing significant phototoxicity.
Beyond these practical applications and experimental paradigms, the relationship between SIMFormer’s reconstruction error and the Cramér–Rao lower bound deserves further investigation to understand the fundamental limits of this approach. Ultimately, the robust and adaptable architecture of SIMFormer shows great potential not only for super-resolution imaging but also for a wide range of other scientific analysis problems where ill-posed inverse problems are a central challenge.
@@
As shown in Extended Data Fig. 1, our encoder is designed to process raw images from SIM and is inspired by Spatiotemporal Masked Autoencoders [Feichtenhofer2022STMAE]. We develop a ViT encoder capable of processing multi-frame SIM images with pre-trained ViT weights (CLIP [Radford2021CLIP]). To speed up training and improve the extraction of fine-grained features, we incorporated ViT-adapter [Chen2023ViTAdapter]. Specifically, our encoder, modified from the ViT-based architecture, converts multi-frame SIM images into multiple 768-dimensional vectors using three-dimensional patch embedding. Simultaneously, the original multi-frame SIM images are processed through a convolutional neural network-based (CNN-based) stem, resulting in an equal number of 768-dimensional vectors for each patch. Both sets are subjected to the same random masking, which masks a portion of the patches. The unmasked patches are combined with learnable position embeddings and frame position embeddings, shuffled, and fed separately into the encoder and adapter layers. These elements interact and merge during forward propagation, facilitating information exchange. The output features of the encoder are then used. A perceptron layer transforms the features of each patch into 512-dimensional vectors. A learnable token is introduced to replace the masked patches. These features are then reorganized into their original order and added with learnable positional embeddings. After this reordering, the features are fed into a decoder consisting of eight transform layers. The output of the decoder is then used by three separate prediction heads.
@@
SIMFormer uses a simple CNN-based prediction module. This module consists of four blocks. Each block is for two-fold upsampling. Within each block, the module performs a double upsampling step, followed by three convolutional operations with filter sizes of $3\times 3$. After each convolution, a gaussian error linear unit activation function is used as the activation function. The prediction module starts with 512 channels and methodically reduces the number of channels in the output of these convolutional layers, halving it in each block. The final stage is a convolutional layer with a $1\times 1$ filter that produces a single channel output.
The emitter prediction subnetwork consists of the prediction module mentioned above, followed by a soft plus activation layer, to achieve non-zero outputs. The background prediction subnetwork incorporates the aforementioned prediction module, coupled with an average pooling layer that features a kernel size of 64. This design assumes that the background exhibits slow intensity variations across the spatial domain.
@@
In SIM, the light pattern consists of only a few modes. For instance, in a typical SIM experiment, the cosine light pattern operates in three directions with nine phases, corresponding to nine distinct modes. This implies that the number of predicted light pattern modes can be reduced. To integrate this prior knowledge into SIMFormer, we employ a low-rank encoding method via an average pooling layer. Specifically, the decoder’s feature output first passes through a Transformer layer, followed by a one-dimensional average pooling operation applied to each feature vector, and then through another Transformer layer before producing the final output.
In predicting the light pattern, a common assumption is that the sum of all light patterns in a set of SIM experiments yields a uniform illumination pattern [Mudry2012SpeckleSIM, Lee2021MetamaterialNanoscopy]. To ensure that the sum equals one and that the intensity of the light patterns remains non-negative, a softmax operation is applied to the output of the prediction module along the channel dimension.
@@
The PSF prediction module is a convolutional neural network that uses a fixed random tensor as input to generate the PSF. Based on the principles of Deep Image Prior [Ulyanov2018DeepImagePrior], the neural network reframes the challenge of estimating the PSF as an optimization task within the neural network. This methodology has been demonstrated to learn the PSF directly from raw microscopy images [Wang2024ImagingMechanism].
@@
We designed several functions and combined them to form a loss function to learn in a self-supervised manner from raw SIM images. These functions can be divided into two categories. The first category measures the difference between the reconstructed SIM images and the raw SIM images. In this category, we introduce the $L_{1}$ distance and multi-scale structural similarity (MS-SSIM). The weighted average of $L_{1}$ distance and MS-SSIM has been proven to be very effective in image restoration tasks [Zhao2016LossFunctions].
In the second category, considering that the intensity-weighted center of the PSF should coincide with the geometric center of the PSF tensor to prevent drift in the prediction results, we designed a center distance function. Additionally, based on optical principles, the intensity of the PSF and light pattern should have spatial continuity, so we introduced a total variation (TV) loss for the PSF and light pattern. The loss function used in the training of SIMFormer is the weighted average of the above functions, as shown below:
$L_{1}$ $=\bigl|\mathrm{SIM}_{\mathrm{pred}}-\mathrm{SIM}_{\mathrm{raw}}\bigr|,$ (1)
$L_{\mathrm{MS\text{-}SSIM}}$ $=1-\mathrm{MS\text{-}SSIM}\bigl(\mathrm{SIM}_{\mathrm{pred}},\mathrm{SIM}_{\mathrm{raw}}\bigr),$ (2)
$\mathrm{Central}$ $=\left|\frac{\sum_{x,y}(x,y)\,\mathrm{PSF}(x,y)}{\sum_{x,y}\mathrm{PSF}(x,y)}-(c_{x},c_{y})\right|,$ (3)
$\mathrm{PSF}_{\mathrm{tv}}$ $=\sum_{x,y}\bigl(\mathrm{PSF}(x+1,y)-\mathrm{PSF}(x,y)\bigr)^{2}+\bigl(\mathrm{PSF}(x,y+1)-\mathrm{PSF}(x,y)\bigr)^{2},$ (4)
$\mathrm{LP}_{\mathrm{tv}}$ $=\sum_{x,y}\bigl(\mathrm{LP}(x+1,y)-\mathrm{LP}(x,y)\bigr)^{2}+\bigl(\mathrm{LP}(x,y+1)-\mathrm{LP}(x,y)\bigr)^{2},$ (5)
$\mathrm{Loss}$ $=\alpha L_{1}+\beta L_{\mathrm{MS\text{-}SSIM}}+\gamma\,\mathrm{Central}+\delta\bigl(\mathrm{PSF}_{\mathrm{tv}}+\mathrm{LP}_{\mathrm{tv}}\bigr).$ (6)
Where $\alpha=0.875$, $\beta=0.125$, $\gamma=0.001$ and $\delta=0.001$.
Training
The training process for SIMFormer begins by loading a pre-trained model, leveraging the benefits of transfer learning since training a ViT from scratch on small datasets may not yield optimal performance. We use the pre-trained CLIP model [Radford2021CLIP] as the initialization parameter for SIMFormer. The initial training is conducted with a batch size of 96 over 70 epochs. The input patch size is set to $3\times 16\times 16$, and the mask ratio is 0.75.
Next, a multi-stage training process is employed to gradually compress the rank of the light pattern representation, adjusting the low-rank coding rank from 512 to 32 (256, 128, 64, and finally 32). In each stage, the mask ratio is 0.25, with each stage running for 30 epochs and batch sizes adjusted to 64. After reducing the low-rank coding rank to 32, the model undergoes an extensive training stage with 2000 epochs, a batch size of 96, and a mask ratio of 0.25 for synthetic data / 0.75 for BioSR. See Supplementary Note 7 for details.
Throughout the process, various training parameters such as the patch size of $3\times 16\times 16$, crop size $80\times 80$, PSF size $49\times 49$, and learning rate of $10^{-4}$ are maintained. The raw images are linearly scaled up by a factor of 3, resulting in threefold enlarged SIM reconstruction images. The unscaled raw images are used when calculating the loss.
@@
We use self-distillation to improve the performance of SIMFormer, a technique where a model is refined by training it to learn from its own predictions (Supplementary Note 8). Specifically, we first train SIMFormer using high signal-to-noise ratio (SNR) SIM images, allowing it to generate high-resolution predictions. These predictions are then used as soft labels to train a new instance of the model SIMFormer+ under conditions of lower photon counts and increased noise. All high signal-to-noise ratio versions of the images in the test set did not appear in the training set of SIMFormer and SIMFormer+. This approach allows SIMFormer+ to inherit the high-resolution capabilities of the original model, while significantly improving its robustness to noise.
@@
Image reconstruction algorithms often produce results with ranges that differ from those of the ground truth images. To address this issue, we apply a commonly used normalization method to ensure a consistent data range. Specifically, percentile normalization is applied to both the ground truth and reconstructed images:
$\mathrm{Norm_{p}(image)}=\frac{\mathrm{image}-\mathrm{Percentile}(\mathrm{image},0.1)}{\mathrm{Percentile}(\mathrm{image},99.9)-\mathrm{Percentile}(\mathrm{image},0.1)}.$
In this equation, we use percentiles of the image to determine the normalization range. $\mathrm{Percentile}(\mathrm{image},0.1)$ represents the 0.1 percentile of the image, and $\mathrm{Percentile}(\mathrm{image},99.9)$ represents the 99.9 percentile of the image.
@@
The resolution is assessed through decorrelation analysis [Descloux2019Decorrelation], which differs from Abbe’s principle by identifying the highest frequency through the examination of local peaks in the decorrelation functions. Details in Supplementary Note 6.
Computational Resource and Speed
In this work, we use the JAX library [Bradbury2018JAX], as detailed in Code File 1. JAX is a widely used open-source deep learning framework known for its efficiency and ease of use in building and deploying models. The model is trained on 8 NVIDIA Tesla A100 80 GB GPUs over a period of $7\times 24$ hours. Inference with the trained model for a single SIM stack $(9,512,512)$ takes approximately 1 second.